Suppose one year ago, Hein Company had inventory in Britain valued at 240,000 pounds. The exchange rate for dollars to pounds was 1£ = 2 U.S. dollars. This year the exchange rate is 1£ = 1.82 U.S. dollars. The inventory in Britain is still valued at 240,000 pounds. What is the gain or loss in inventory value in U.S. dollars as a result of the change in exchange rates?
1
Expert's answer
2012-06-19T10:41:14-0400
Previous value of inventories in dollars: 240000*2=480000$.



Current value of inventories in dollars: 240000*1.82=436800$



As we see, company experienced a loss of 480000-436800=43200$.